OpenStudy (anonymous):

what expression repersents the volume of a cube that is 2x on a edge

Parth (parthkohli):

\( \color{Black}{\Rightarrow edge^3 }\)

OpenStudy (kropot72):

What is the area of one face of the cube?

OpenStudy (anonymous):

It just gives me answer and though are 6x 6x^3 8x^3 24x^2

OpenStudy (kropot72):

The area of one face is (2x) * (2x) = 4 * x^2 Now multiply this result by 2x and you have the answer.

OpenStudy (anonymous):

What's the volume of a regular pyramid if the length is 3 and the width is 2

OpenStudy (kropot72):

Fot the cube \[4\times x ^{2}\times 2x=?\]

OpenStudy (anonymous):

A pentagon prism has a volume of 448 cm^3 If the hieght of the prism is 7cm what is the area of the base?